OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. x + 2y = 12 y = x + 3

OpenStudy (anonymous):

(2,5)

OpenStudy (anonymous):

2x + 3y = 13 x = 2

OpenStudy (anonymous):

x+2(x+3) = 12 x+2x+6 = 12 3x = 12-6 x = 6/3 x = 2 y = 2+3 y = 5

OpenStudy (anonymous):

2x + 3y = 13 x = 2 3y=13-4 3y=9 y=3